### v1

@@ 3.3 Self-supervised Spike-guided Deblurring Model
To improve them, we further build a knowledge distillation framework building upon the existing processing pipeline, which provides the reconstructed sequence as pseudo-labels for the training of the student network LDN as shown in Fig. 3. LDN adheres to a similar input and output pattern as [8], i.e., taking the blurry input $\mathbf{B}$ and the short spike stream $\mathcal{S}_{\mathcal{T}^{\prime}}$ centered around moment $t$ as inputs, with the output being the reconstructed sharp image $\mathbf{L}_{ldn}(t)$. The mapping relationship of the LDN can be mathematically formulated as follows:
$\mathbf{L}_{ldn}(t)=\text{LDN}(\mathbf{B},\mathcal{S}_{\mathcal{T}^{\prime}};\Theta_{3}),$ (11)
where $\Theta_{3}$ represents the parameters of the LDN. The LDN is designed according to the encoder-fusion-decoder paradigm [23, 4, 8]. Full details regarding its structure are available in the supplementary materials.
In real-world scenarios where ground truth images are unavailable for supervision, we employ the previously described processing pipeline as the teacher model. To avoid the scenario where the LDN exactly replicates the mapping of the teacher model, we construct the teacher loss $\mathcal{L}_{tea}$ based on the LPIPS [29] and further introduce the blur reconstruction loss. The reblur loss $\mathcal{L}_{reblur}$ measures the difference between the blurry input $\mathbf{B}$ and the re-synthesized blurry image ${\widetilde{\mathbf{B}}}$, satisfying:
$\widetilde{\mathbf{B}}=\displaystyle\sum_{m=0}^{M-1}\mathbf{L}_{ldn}(t_{m}),$ (12)
where $\mathbf{L}_{ldn}(t_{m})$ represents the $m$-th recovered image within the exposure period $\mathcal{T}$ and $M$ is the total number of reconstructed images. To sum up, the final loss function for our knowledge distillation framework is formulated as follows:
$\mathcal{L}$ $=\mathcal{L}_{tea}+\lambda\cdot\mathcal{L}_{reblur}$ (13)
$=\sum_{m=0}^{M-1}\mathcal{L}_{LPIPS}(\mathbf{L}_{bsn}^{\uparrow}(t_{m}),\mathbf{L}_{ldn}(t_{m}))+\lambda\cdot\mathcal{L}_{MSE}(\widetilde{\mathbf{B}},\mathbf{B}),$ (14)
where $\mathbf{L}_{bsn}^{\uparrow}(t_{m})$ represents the denoising result after being upscaled by the SR network, and $\lambda$ is a weighting parameter.
@@ 4.1 Dataset
Synthetic Data. For quantitative and qualitative analysis of our spike-guided motion deblurring task, we construct the synthetic dataset based on the widely employed motion deblur dataset GOPRO [17] with the resolution of $1280\times 720$. Initially, we utilize the interpolation algorithm XVFI [21] to augment the video frame by interpolating additional $7$ frames between each pair of consecutive sharp images. To generate the spike stream that mimics reality closely, we downsample the interpolated video to the resolution of $320\times 180$ and simulate the spike stream based on the spike simulator [34] modeled according to the spike generation physical model. To replicate real-world motion blur, we synthesize each blurry input by averaging $97$ frames from interpolated video sequences.
Real-world Data. To comprehensively assess the deblurring performance of our S-SDM in real-world scenarios, we construct an RGB-Spike binocular system and propose the first Real-world Spike-guided Blur (RSB) dataset in this community. This system consists of a fixed-exposure RGB camera (Basler acA1920-150uc) and a spike camera [10], enabling us to capture the blurry image and the corresponding spike stream simultaneously. Further details about our RSB and the spatial-temporal calibration for our binocular system are provided in the supplementary materials.
Table 1: Quantitative comparison of the single frame and sequence reconstruction task on the GOPRO dataset, where ‘GT’ indicates that this method requires ground truth for training and ‘Single Frame’ refers to the middle frame of the restored sequence.
Methods
Inputs
GT
Single Frame Sequence
Params
RGB
Spike
PSNR $\uparrow$
SSIM $\uparrow$
PSNR $\uparrow$
SSIM $\uparrow$
LEVS [12]
$\times$
25.400
0.727
21.155
0.601
4.97M
Motion-ETR [32]
$\times$
27.821
0.763
21.955
0.610
6.55M
BiT [38]
$\times$
24.442
0.724
23.644
0.698
11.3M
SpkDeblurNet [8]
32.108 0.873 30.712 0.821
13.4M
S-SDM (Ours)
$\times$
27.928 0.786 26.893 0.757 0.234M
@@ 4.2 Experimental Results
We conduct both quantitative and qualitative comparisons of our S-SDM against state-of-the-art (SOTA) methods, including frame-based Motion-ETR [32], LEVS [12], video-based BiT [38] and spike-based SpkDeblurNet [8] on the GOPRO and RSB datasets. Since SpkDeblurNet does not fully consider the real spatial-resolution gap between two modalities in real-world scenarios, we modify its spike down-sampling module to perform convolution at the original size and retrain it on our synthetic dataset for a fair comparison.
GOPRO Dataset. We reconstruct $7$ images from one blurry input for sequence restoration evaluation [31] as detailed in Tab. 1. Given that the principal contribution of the S-SDM is self-supervised learning, it is foreseeable that our method might be slightly inferior to the supervised method SpkDeblurNet on the GOPRO synthetic dataset. Revisiting the mentioned motion ambiguity problem, we observe that frame-based methods LEVS and Motion-ETR suffer significant performance degradation in transitioning from single frame restoration to sequence reconstruction, with decreases of 4.425 dB $\downarrow$ and 5.866 dB $\downarrow$ in PSNR respectively. In contrast, our method alleviates this issue benefiting from the incorporation of the high temporal-resolution spike stream.
@@
RSB Dataset. We further present visualizations of single frame and sequence reconstruction comparisons on the real-world RSB dataset. For the single frame restoration qualitative experiments as depicted in Fig. 4, both frame-based and video-based approaches fail to replicate fine textures and detailed elements present in the blurry input. While SpkDeblurNet is capable of recovering structural details, it is deteriorated by significant noise and color distortion under conditions with lower spike density than those simulated in the GOPRO dataset. Similarly, in scenarios of higher spike stream densities, the restoration of SpkDeblurNet tends to exhibit over-exposure, resulting in brightness inconsistency compared to the blurry input. In terms of high-frequency details, other methods struggle to eliminate the laser ghost as evident in the earphone scene of Fig. 4.
In the qualitative experiments focusing on sequence restoration as in Fig. 5, LEVS fails to analyze motion information during the exposure, resulting in uniform outcomes across each image in the sequence. While Motion-ETR and BiT appear to trace the moving pattern of the calibration board correctly, the texture restoration is severely destroyed. Benefiting from the incorporation of the spike stream, though SpkDeblurNet correctly recovers the trajectory, it encounters deterioration problems similar to those observed in Fig. 5. Our method tackles the aforementioned challenges in a self-supervised manner, which facilitates re-training on the real-blur dataset, ensuring that the restored sequence is aligned with the real-captured blurry input and spike stream. More comparative experiments and analyses are accessible in the supplementary material.
Table 2: Ablation study for evaluating the modules of S-SDM.
Table 3: Ablation study for evaluating the effectiveness of the distillation framework.
ID
BSN
SR
LDN
@@
I-1
$\times$
22.943
0.461
I-2
$\checkmark$
24.490
0.645
I-3
$\checkmark$
$\checkmark$
26.144
0.708
I-4
$\checkmark$
$\checkmark$
$\checkmark$
27.928 0.786
ID
$\mathcal{L}_{tea}$
$\mathcal{L}_{reblur}$
$\lambda$
@@
II-1
$\times$
$\checkmark$
10
23.102
0.441
II-2
$\checkmark$
$\times$
26.713
0.723
II-3
$\checkmark$
$\checkmark$
10
27.380
0.762
II-4
$\checkmark$
$\checkmark$
50
27.804
0.779
II-5
$\checkmark$
$\checkmark$
70
27.825
0.780
II-6
$\checkmark$
$\checkmark$
100
27.928 0.786
Table 3: Ablation study for evaluating the effectiveness of the distillation framework.
4.3 Ablation Study
In this subsection, we perform ablation experiments on the GOPRO and RSB datasets to evaluate the performance of each module within S-SDM, as well as the overall effectiveness of our distillation learning framework.
Modules Cascading. Building upon the SDM, we sequentially cascade the denoising network, the SR network, and the LDN to evaluate their respective effectiveness. The quantitative ablation experiments are presented in Tab. 3. Given that the spatial resolution of the spike camera is one-quarter of that of conventional cameras, we employ nearest neighbor interpolation as a substitute for the SR network in experiments I-1 and I-2 to align the spatial resolution of two modalities.
Qualitative comparisons based on the GOPRO dataset are illustrated in Fig. 6. These comparisons reveal that although the SDM is effective in removing motion blur, it struggles with significant noise and detail loss, attributed to the inherent spike noise and the relatively low spatial resolution of the spike camera. While the BSN mitigates noise and the SR network improves image resolution explicitly, the LDN trained via distillation learning further refines these enhancements, enabling the recognition of intricate textural features in the images, such as the license plate and the door number shown in Fig. 6.
[…]

### v4

@@ 3.3.2 Knowledge Distillation Framework
To improve them, we further build a knowledge distillation framework building upon the existing processing pipeline. This pipeline serves as the teacher model, providing the reconstructed sequence as pseudo-labels for the training of the student model LDN, as illustrated in Fig. 2.
Lightweight Deblur Network. LDN adheres to a similar input and output pattern as previous research [8], i.e., taking the blurry input $\mathbf{B}$ and the short spike stream $\mathcal{S}_{\mathcal{T}^{\prime}}$ centered around moment $t$ as inputs, with the output being the reconstructed sharp image $\mathbf{L}_{\text{ldn}}(t)$, mathematically formulated as follows:
$\mathbf{L}_{\text{ldn}}(t)=\text{LDN}(\mathbf{B},\mathcal{S}_{\mathcal{T}^{\prime}};\Theta_{3}),$ (11)
Full details regarding the LDN structure are available in the supplementary materials.
To avoid the scenario where the LDN exactly replicates the mapping of the teacher model, we design the teacher loss $\mathcal{L}_{\text{tea}}$ based on the LPIPS [34] loss and further introduce the blur reconstruction loss. The reblur loss $\mathcal{L}_{\text{reblur}}$ measures the difference between the blurry input $\mathbf{B}$ and the re-synthesized blurry image ${\widetilde{\mathbf{B}}}$, satisfying:
$\widetilde{\mathbf{B}}=\frac{1}{M}\displaystyle\sum_{m=1}^{M}\mathbf{L}_{\text{ldn}}(t_{m}),$ (12)
where $\mathbf{L}_{\text{ldn}}(t_{m})$ represents the $m$-th recovered image within the exposure period $\mathcal{T}$ and $M$ is the total number of reconstructed images. Finally, we sum up two loss functions with the weighting parameter $\lambda$, and the final loss function is formulated as follows:
$\mathcal{L}$ $=\mathcal{L}_{\text{tea}}+\lambda\cdot\mathcal{L}_{\text{reblur}}$ (13)
$=\sum_{m=1}^{M}\mathcal{L}_{\text{LPIPS}}(\mathbf{L}_{\text{bsn}}^{\uparrow}(t_{m}),\mathbf{L}_{\text{ldn}}(t_{m}))+\lambda\cdot\mathcal{L}_{\text{MSE}}(\widetilde{\mathbf{B}},\mathbf{B}).$ (14)
@@ 4.1 Dataset
Synthetic Data. For quantitative analysis of our spike-guided motion deblurring task, we construct the synthetic dataset based on the widely employed GOPRO [18] dataset. We initially utilize the interpolation algorithm XVFI [23] to augment the video frame by interpolating additional $7$ frames between each pair of consecutive sharp images. To generate the spike stream that mimics reality closely, we downsample the interpolated video to the resolution of $320\times 180$ and simulate the spike stream based on the spike simulator [41]. To replicate real-world motion blur, we synthesize each blurry input by averaging $97$ frames from interpolated video sequences.
Real-world Data. We construct an RGB-Spike binocular system and propose the first Real-world Spike-guided Blur dataset (RSB) in this community. This system consists of a fixed-exposure RGB camera (Basler acA1920-150uc) and a spike camera [11], enabling us to capture the blurry image and the corresponding spike stream simultaneously. Further details about our RSB and the spatial-temporal calibration for our binocular system are provided in the supplementary materials.
Table 1: Quantitative comparison of the sequence reconstruction task on the GOPRO dataset.
{tblr}
width = colspec = X[5,l]X[1,c]X[0.05,c]X[1.3,c]X[1.3,c]X[0.05,c]X[1.3,c]X[1.3,c]X[0.05,c]X[1.3,c]X[1.3,c]X[1,c], cell11 = r=2, cell12 = r=2, cell14 = c=2, cell17 = c=2, cell110 = c=2, cell112 = r=2, hline3,10 = -, hline2 = 4-5, hline2 = 7-8, hline2 = 10-11, hline1 = 1-122pt, hlineZ = 1-122pt, Methods & Spike $V_{th}$=1 $V_{th}$=2 $V_{th}$=4 Params
PSNR SSIM PSNR SSIM PSNR SSIM
LEVS[13] $\times$ 21.155 0.601 21.155 0.601 21.155 0.601 4.97M
Motion-ETR[39] $\times$ 21.955 0.610 21.955 0.610 21.955 0.610 6.55M
BiT[49] $\times$ 23.644 0.698 23.644 0.698 23.644 0.698 11.3M
TRMD[4]+DASR[27] ✓ 27.323 0.784 21.198 0.601 18.567 0.523 19.3M
RED[30]+DASR[27] ✓ 24.456 0.741 23.178 0.674 21.942 0.608 9.76M
REFID[26]+DASR[27] ✓ 28.124 0.819 15.288 0.339 13.623 0.274 15.9M
SpkDeblurNet[8]28.307 0.834 14.406 0.299 11.621 0.202 13.4M
S-SDM (Ours) ✓ 26.893 0.757 26.367 0.740 25.433 0.699 0.23M
@@ 4.2 Experimental Results
We conduct both quantitative and qualitative comparisons of our S-SDM against state-of-the-art (SOTA) motion deblurring methods, including frame-based Motion-ETR [39], LEVS [13], video-based BiT [49], event-based TRMD [4], REFID [26], RED [30] and spike-based SpkDeblurNet [8] on the GOPRO and RSB datasets. For event-based methods, we replace the event stream with the spike stream and adopt the same input representation in these methods [30, 26, 4]. We further cascade the image super-resolution technique DASR [27] as in [37] for the deblurred sequence to overcome the modality inconsistency which is not considered in these methods. We reconstruct $7$ images from one blurry input for sequence restoration evaluation [37] as listed in Tab. 1.
Results on GOPRO. To simulate the spike density domain gap as depicted in Fig. 1, we train all supervised spike-based deblurring methods on the GOPRO dataset under spike threshold [41] $V_{th}=1$ and evaluate them on datasets with spike thresholds $V_{th}=1,2,4$. Quantitative comparison results are listed in Tab. 1. Given that the principal contribution of our S-SDM is self-supervised learning, it is foreseeable that our method might be slightly inferior to the supervised methods SpkdDeblurNet on the dataset with $V_{th}=1$. While these supervised methods deteriorate on datasets with $V_{th}\neq 1$, our method achieves great generalization benefiting from the self-supervision design. Besides, our method achieves better restoration performance than the self-supervised method RED due to our consideration of the spatial-resolution mismatch between two modalities and the designed teacher loss, which imposes a stronger constraint than the optical loss in RED.
@@
Results on RSB. We further present visualizations of single frame and sequence reconstruction comparisons on the real-world RSB dataset as depicted in Fig. 3 and 4. Both frame-based and video-based approaches fail to replicate fine textures and detailed elements present in the blurry input. While SpkDeblurNet is capable of recovering structural details and the motion trajectory, it is deteriorated by significant noise and color distortion under conditions with lower spike density than those simulated in the GOPRO dataset. Similarly, in scenarios of higher spike stream densities, the restoration of SpkDeblurNet tends to exhibit over-exposure, resulting in brightness inconsistency compared to the blurry input. Our method addresses these challenges by finetuning on the RSB dataset, ensuring that the restored sequence aligns with the real-captured blurry input and spike stream. More comparative experiments and analyses are accessible in the supplementary material.
4.3 Ablation Study
We perform ablation experiments on the GOPRO and RSB datasets to evaluate the performance of each module within S-SDM, as well as the overall effectiveness of our distillation learning framework. In this subsection, we evaluate performance using the single middle frame for simplicity.
Modules Cascading. Building upon the SDM, we sequentially cascade the BSN, the SR, and the LDN to evaluate their respective effectiveness, with quantitative results on GOPRO presented in Tab. 3. We employ bilinear interpolation as a substitute for the SR network in experiments I-1 and I-2 to align the spatial resolution of two modalities.
Qualitative ablation experiments are illustrated in Fig. 6. These comparisons reveal that while the SDM effectively removes motion blur, it struggles with significant noise and detail loss due to the spike noise and the low resolution of the spike stream. While the BSN mitigates noise and the SR network improves spatial resolution explicitly, the LDN trained via distillation learning further refines these enhancements, enabling the recognition of intricate textural features in the images, such as the license plate and the door number shown in Fig. 6.
Figure 5: Modules cascading comparisons on GOPRO. Experiments ID can be viewed on Tab. 3.
Figure 6: Distillation learning comparisons on RSB. Experiments ID can be viewed on Tab. 3
Table 2: Modules cascading ablation on GOPRO.
@@ 4.3 Ablation Study
I-1 $\times$ 23.012 0.486
I-2 $\checkmark$ 24.634 0.661
I-3 $\checkmark$ $\checkmark$ 26.144 0.708
I-4 $\checkmark$ $\checkmark$ $\checkmark$ 27.928 0.786
Table 3: Distillation learning ablation on GOPRO.
@@
II-1 $\times$ $\checkmark$ 10 23.102 0.441
II-2 $\checkmark$ $\times$ / 26.563 0.723
II-3 $\checkmark$ $\checkmark$ 10 27.345 0.762
II-4 $\checkmark$ $\checkmark$ 50 27.742 0.778
II-5 $\checkmark$ $\checkmark$ 100 27.928 0.786
II-6 $\checkmark$ $\checkmark$ 200 27.620 0.783
Distillation Learning. We focus on analyzing the contribution of the teacher loss $\mathcal{L}_{\text{tea}}$ and the reblur loss $\mathcal{L}_{\text{reblur}}$ within the distillation framework, with quantitative results listed in Tab. 3. Without the teacher loss $\mathcal{L}_{\text{tea}}$, the LDN tends toward learning identity mapping from the blurry input. While under the guidance of the teacher model, the reblur loss $\mathcal{L}_{\text{reblur}}$ not only enforces motion consistency in the reconstructed sequence but also enriches the LDN with high-resolution details from the non-blurry regions of the input, thus improving the performance on GOPRO as listed in Tab. 3.
We further apply the LDN trained on GOPRO to the real-world dataset RSB, with qualitative visualization illustrated in Fig. 6. As observed in the figure, the absence of the reblur loss $\mathcal{L}_{\text{reblur}}$ leads to significant noise in the recovered image, which predominantly arises from the disparity in spike density and generation mechanism between the simulated and real-captured spike stream. This discrepancy causes the LDN to overestimate the spike number, resulting in black holes in regions with lower spike density than simulated, which reflects the drawback inherent in the supervised learning strategies as discussed in Sec. 1. Increasing the weight of the reblur loss $\mathcal{L}_{\text{reblur}}$ allows the LDN to incorporate more information from the blurry input, thereby mitigating this issue. However, this adjustment also leads to the presence of blurry edges. We follow the parameters set in Experiment II-6 and retrain the LDN on the RSB dataset (referred to as Experiment II-7). The retrained LDN effectively recovers the sharp edge of the calibration board and suppresses the spike noise in the background, validating the feasibility of our self-supervised framework in real-world scenarios.
5 Conclusion
In this paper, we introduce a novel self-supervised spike-guided motion deblurring framework S-SDM, which reconstructs sequences of sharp images from real-world blurry inputs with the spike stream. Additionally, we construct an RGB-Spike binocular system and propose the first spatially-temporally calibrated real-world dataset RSB in this community. Quantitative and qualitative experiments validate the superior generalization capabilities of our proposed S-SDM.
Limitation. The limitation of our S-SDM lies in its dependence on strict spatial-temporal calibration. Misalignment will lead to color shifts and quality degradation in the deblurred sequence.
Appendix A Appendix
@@
[width=autoplay, loop]3imgs/Sup/middle/middle_029
@@
[width=autoplay, loop]3imgs/Sup/high/high_029
Figure 7: Video comparison of our method against the BiT and SpkDeblurNet on the RSB dataset under different luminance conditions, with the static visualization displayed in Fig. 12. It is recommended to view the pdf using Acrobat PDF reader and the gif demonstration with a higher frame rate is available in the supplementary zip file.
This supplementary material provides a comprehensive elaboration on the methodologies and experiments in this paper. It is organized into four distinct sections: Theory Analysis in Sec. A.1, Network Settings in Sec. A.2, RSB Dataset in Sec. A.3, Experimental Details in Sec. A.4 and Additional Figure in Sec. A.5.
A.1 Theory Analysis
We define $k_{1}^{\mathbf{B}}$ and $k_{2}^{\mathbf{B}}$ as the ratios of the red channel to the green and blue channels in the blurry input $\mathbf{B}$ respectively, i.e.,
$k_{1}^{\mathbf{B}}$ $=\mathbf{B}_{r}/\mathbf{B}_{gre},$ (15)
$k_{2}^{\mathbf{B}}$ $=\mathbf{B}_{r}/\mathbf{B}_{b}.$ (16)
Since the gray image is the weighted sum of RGB channels, the ratio of the gray to the red image $\alpha_{r}^{\mathbf{B}}$ is formulated as:
$\alpha_{r}^{\mathbf{B}}$ $=\mathbf{B}_{g}/\mathbf{B}_{r}$ (17)
$=(w_{r}\cdot\mathbf{B}_{r}+w_{gre}\cdot\mathbf{B}_{gre}+w_{b}\cdot\mathbf{B}_{b})/\mathbf{B}_{r}$ (18)
$=w_{r}+w_{gre}/k_{1}^{\mathbf{B}}+w_{b}/k_{2}^{\mathbf{B}},$ (19)
where $w_{c}$ denote the weight of channel $c\in\{r,gre,b\}$ respectively.
Similarly, we define $k_{1}^{\mathbf{E}}(t,\mathcal{T}^{\prime})$ and $k_{2}^{\mathbf{E}}(t,\mathcal{T}^{\prime})$ as the fractions of the red channel relative to the green and blue channels in the short-exposure image $\mathbf{E}(t,\mathcal{T}^{\prime})$, resulting in:
$\alpha_{r}^{\mathbf{E}}(t,\mathcal{T}^{\prime})=w_{r}+w_{gre}/k_{1}^{\mathbf{E}}(t,\mathcal{T}^{\prime})+w_{b}/k_{2}^{\mathbf{E}}(t,\mathcal{T}^{\prime}).$ (20)
Given the observation that the color information of adjacent pixels often exhibits similarity under the premise of minor motion amplitude, we postulate that the colors of the blurry input $\mathbf{B}$ and the short-exposure image $\mathbf{E}(t,\mathcal{T}^{\prime})$ are identical. This assumption implies that the intensity proportion among RGB channels in two images is approximately equivalent, i.e., $k_{1}^{\mathbf{B}}\approx k_{1}^{\mathbf{E}}(t,\mathcal{T}^{\prime})$ and $k_{2}^{\mathbf{B}}\approx k_{2}^{\mathbf{E}}(t,\mathcal{T}^{\prime})$. Drawing from Eq. 19 and Eq. 20, We further deduce the following relation:
$\alpha_{r}^{\mathbf{B}}\approx\alpha_{r}^{\mathbf{E}}(t,\mathcal{T}^{\prime}),$ (21)
which can be readily generalized across channels, leading to $\alpha_{c}^{\mathbf{B}}\approx\alpha_{c}^{\mathbf{E}}(t,\mathcal{T}^{\prime})$.
Figure 8: Network diagrams of the BSN (a) and our designed LDN (b).
A.2 Network Settings
A.2.1 Blind Spot Network
We construct our BSN based on the blind-spot strategy outlined in [14]. We first rotate the input image four times, then concatenate them into the U-Net [21] structure like denoising network. To prevent direct mapping from the input pixel to the output pixel, a single-pixel offset strategy is employed in the convolutional kernel to separate its receptive field from the central pixel. Finally, the denoising output is obtained by merging the results of four branches via a $1\times 1$ convolution, as shown in Fig. 8(a).
A.2.2 Lightweight Deblur Network
The network structure is depicted in Fig. 8(b), where the encoder for blurry images consists of two layers of down-sampling convolutions to align the spatial resolution of two modalities. In contrast to the intricate cross-attention mechanism outlined in [4, 8], we implement the fusion of two modalities by the simple operation of $\text{Concat}(\cdot)$, with the cascaded CBAM [29] and residual blocks for further feature fusion.
A.3 RSB Dataset
We detail the construction of our RGB-Spike hybrid system as illustrated in Fig. 10. The system comprises a Spike Camera-001T-Gen2 with a resolution of 400 $\times$ 250 pixels, paired with a Basler acA1920-150uc RGB Camera, offering a higher resolution of 1920 $\times$ 1200 pixels.
Figure 9: RGB-Spike camera system.
Figure 10: Calibration result.
To achieve the spatial calibration of two cameras, we perform simultaneous captures of the calibration board using the RGB-Spike system as shown in Fig. 10. The RGB images are cropped to a resolution of 1600 $\times$ 1000, aligning them to be fourfold the resolution of the spike camera. We convert the cropped RGB image to grayscale and take the TFP [50] image reconstruction result as the reference from the spike camera. We utilize the MATLAB calibration toolbox to implement the calibration process, with the results detailed in Fig. 10.
Table 4: Quantitative comparison of the single frame task on the RSB dataset, where non-reference metric LIQE ranging from 1 to 5 is employed. LIQE is a positive metric denoted as $\uparrow$ where higher scores reflect better performance.
Methods Board-L Board-M Board-H Face Earphone Average $\uparrow$
LEVS [13] 1.0010 1.0342 1.0400 1.0029 1.1382 1.0433
Motion-ETR [39] 1.0153 1.0064 1.0279 1.0023 1.1724 1.0449
BiT [49] 1.0002 1.0069 1.0473 1.0138 1.6155 1.1367
SpkDeblurNet [8] 1.8493 1.3232 2.2698 1.2210 2.6606 1.8648
S-SDM (Ours) 2.0955 1.3872 2.4208 1.3814 2.7087 2.1987
Figure 11: Qualitative comparisons for single-frame restoration on the RSB dataset are illustrated, where “Board-L”, “Board-M”, and “Board-H” represent the board captured under low, middle and high lighting conditions.
Our RSB dataset contains 10 video sequences under different conditions, captured under varied conditions including scene brightness levels (e.g., Low, Middle, and High light) and motion patterns (e.g., camera shake and object motion), which introduce different types of motion blur. Besides, the RSB dataset comprises a large amount of blur-spike pairs with each blurry input corresponding to 400 spike frames.
A.4 Experimental Details
A.4.1 Comparison
To assess the performance of our method on the GOPRO dataset, we utilize the Peak Signal-to-Noise Ratio (PSNR) and the Structural Similarity Index (SSIM) as the quantitative metrics, which are commonly used in motion deblurring tasks. In real-world datasets, where obtaining ground truth sharp sequence is challenging, we opt for the non-reference image quality assessment method Language-Image Quality Evaluator (LIQE) [35] as a reference. By assessing visual quality through the computation of joint probabilities from visual-textual embeddings, LIQE adeptly identifies the clarity and blurriness of images independently of the ground truth, making it ideally suited for our task.
Comparison on the RSB dataset. Qualitative and quantitative experiments of the single frame restoration task on the RSB dataset are shown in Fig. 11 and 4 respectively. The visual results coupled with the LIQE metrics demonstrate that our method outperforms other methods in handling the real-world RSB dataset. While the supervised SpkDeblurNet encounters substantial noise and overexposure challenges in both low-light and high-light environments, our approach demonstrates superior restoration performance, which is attributed to the designed self-supervised framework.
Blur Refer to caption
SpkDeblurNet [8] Refer to caption Ours
(a) Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption(b) Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption(c)Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption(d) Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionOursRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Blur Refer to caption
SpkDeblurNet [8] Refer to caption Ours
(a) Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption(b) Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption(c)Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption(d) Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionOursRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 12: Qualitative comparison for the sequence reconstruction on the RSB dataset. (a),(b),(c),(d) denote results of “LEVS", “Moiton-ETR" , “BiT" and “SpkDeblurNet" respectively. The upper panel depicts a real-world scene with a lower spike density than the simulation, whereas the lower image exhibits a higher spike density.
Moreover, as depicted in Fig. 12 and Fig. 7, S-SDM exhibits outstanding performance in precisely retrieving luminance information and texture details while ensuring the motion consistency of reconstructed sequences. In contrast, the video reconstruction of the BiT suffers from poor image quality and shows inadequate sequence continuity in the restored sequence. Besides, SpkDeblurNet encounters issues with color distortion, brightness inconsistency, and inaccurate texture restoration owing to the domain gap between synthetic and real-world datasets. These observations further highlight the superior performance of our S-SDM in real-world scenarios.
Comparison on the GOPRO dataset. We further provide the visual comparison of our method against other SOTA methods on the GOPRO dataset as shown in Fig. 13, 14, 15 and 16. It
Figure 13: Comparison of our S-SDM against other methods on GOPRO with $V_{th}=1,2,4$.
Figure 14: Comparison of our S-SDM against other methods on GOPRO with $V_{th}=1,2,4$.
Figure 15: Comparison of our S-SDM against other methods on GOPRO with $V_{th}=1,2,4$.
Figure 16: Comparison of our S-SDM against other methods on GOPRO with $V_{th}=1,2,4$.
A.4.2 Implementation
While the training of our S-SDM involves training in BSN, EDSR, and LDN, respectively, only LDN is utilized to obtain the deblurred sequence during the evaluation stage. To augment the dataset and accelerate the training process, we randomly crop $512\times 512$ image from each blurry frame, along with the $128\times 128$ spike stream. We use PyTorch to build and train our S-SDM using an NVIDIA GeForce GTX 4090 GPU and AMD EPYC 7742 64-Core Processor. The training of our LDN consumes about 4 hours on the GOPRO. During the testing phase, we feed the entire image and the spike stream into the network to assess performance.
We complete the training of BSN on the GOPRO dataset, employing an initial learning rate of $3e^{-4}$ and spanning 1000 epochs. The training uses the Adam optimizer with a cosine scheduler and sets the batch size to 8 for each epoch. Adopting the same settings as BSN, EDSR is trained on the blur-spike paired data for 70 epochs. Subsequently, LDN undergoes the training of 100 epochs with the learning rate adjusted to $1e^{-3}$.